the length of a rectangle is 31 centimeters less than five times its width. the area is 72 square centimeters. find the dimensions of the rectangle.
Answer by mananth(16946)

let the width be x
length will be 5x-31
Area = L*W
72 = x(5x-31)
72=5x^2-31x
5x^2-31x-72=0
.
x1,x2 the roots of the equation will be
x1= 31+sqrt(-31^2-4*5*-72) / 2*5
x1= 31 +sqrt(-31^2 +1440) /10
x1= 8
x2 = -1.8
x2 is negative
So width = 8 cms
Then length will be 9cms